The mass moment of inertia I of a homogeneous sphere about its diameter is I = (2/5)m R^2, where m and R are its mass and radius, respectively. Find the dimension of I in items of the absolute {
1
Expert's answer
2018-04-12T09:54:11-0400
I = (2/5)m R^2
[I ]= [(2/5)m R^2 ]
[I ]= [(2/5)][m ][ R^2 ]=1∙M∙L^2=ML^2.